OpenStudy (anonymous):

log4(x)=1/2 help please :C

OpenStudy (anonymous):

\[\log_4(x) = \frac{1}{2}\] \[4^{log_4(x)} = 4^{\frac{1}{2}}\] \[x = 4^{\frac{1}{2}} = 2\]